An electric generator delivers a load current of 24A at a terminal voltage of 250v. The generator is driven by a motor,whose output Power is 7.5kw. What is the efficiency of the generator?
1
Expert's answer
2020-12-10T11:09:06-0500

The power given into the load is:


"P_{actual} = UI = 250V\cdot 24A = 6000W = 6.0kW"

The power obtained from the motor is:


"P_{obtained} = 7.5kW"

By definition, the efficiency of the generator is


"\eta = \dfrac{P_{actual}}{P_{obtained}}\times 100\% = \dfrac{6.0kW}{7.5kW}\times 100\% =80\%"

Answer. 80%.